### v1

# Input Reduction Enhanced LLM-based Program Repair

Abstract.
Large Language Models (LLMs) have shown great potential in Automated Program Repair (APR). Test inputs, being crucial for reasoning the root cause of failures, are always included in the prompt for LLM-based APR. Unfortunately, LLMs struggle to retain key information in long prompts. When the test inputs are extensive in the prompt, this may trigger the “lost-in-the-middle” issue, compromising repair performance. To address this, we propose ReduceFix, an LLM-based APR approach with a built-in component that automatically reduces test inputs while retaining their failure-inducing behavior. ReduceFix prompts an LLM to generate a reducer that minimizes failure-inducing test inputs without human effort, and then feeds the reduced failure-inducing inputs to guide patch generation.
For targeted evaluation, we constructed LFTBench, the first long-input APR benchmark with 200 real bugs from 20 programming tasks, each paired with a failure‑inducing input whose median size is 1 MB. On this benchmark, ReduceFix shrinks inputs by 89.1% on average and improves overall pass@10 by up to 53.8% relative to a prompt that includes the original test, and by 17.6% compared with omitting the test entirely. Adding the same reduction step to ChatRepair increases its fix rate by 21.3% without other changes. Ablation studies further highlight the impact of input length and compressed failure information on repair success. These results underscore that automatically reducing failing inputs is a practical and powerful complement to LLM-based APR, significantly improving its scalability and effectiveness.
copyright: none
Figure 1. Motivating example (ABC376C): input reduction shrinks a ¿3MB failure-inducing test to three critical lines, enabling the LLM to generate the accurate patch.
1. Introduction
APR aims to automatically generate bug-fixing patches for software defects, thereby reducing the manual effort required for debugging (Yang et al., 2025a; Le Goues et al., 2019; Xu et al., 2024). Recently, LLMs have been applied to APR with promising results (Xia and Zhang, 2022; Luo et al., 2025). Many recent APR systems enhance patch generation by including test inputs in the prompt, among which the test suite plays a particularly important role (Kong et al., 2024; Tang et al., 2024; Yang et al., 2024). It provides a concrete example of how the program succeeds or fails, helping the LLM focus on the underlying issue and produce a correct fix. This strategy has shown strong results in recent systems such as ChatRepair (Xia and Zhang, 2024). Existing APR studies typically evaluate on benchmarks such as Defects4J (Just et al., 2014) and HumanEval-Java (Jiang et al., 2023), where test inputs are short and rarely exceed a few hundred characters. However, when the test input becomes too long, it is difficult to pinpoint the root cause of the error, known as the “lost-in-the-middle” phenomenon (Liu et al., 2023; Pham et al., 2015), where information buried in a long prompt receives little attention and overall task performance drops (Yang et al., 2024; Tian et al., 2023). Therefore, an automatic test input reduction step becomes essential before repair.
Existing works on test input reduction rely heavily on human effort, which can be divided into two main categories.
Syntax-based approaches, such as HDD (Misherghi and Su, 2006) and Perses (Sun et al., 2018), are built on the classical ddmin algorithm (Zeller and Hildebrandt, 2002) but incorporate grammar or tree structure to ensure syntactic validity during reduction. However, adapting these APR tools to new tasks requires designing a new grammar and tuning heuristics for each input format, which is both time-consuming and error-prone. Other techniques, such as ddSMT (Niemetz and Biere, 2013) and J-Reduce (Kalhauge and Palsberg, 2021), target specific domains, but still require domain knowledge and significant manual effort to implement and are not reusable across different formats. In LLM-based APR, the input format varies widely across tasks, often involving different formats ranging from plain text to structured JSON or domain-specific encoding (Liu et al., 2020). As a result, relying on hand-crafted reducers is not scalable and makes automation difficult.
These observations uncover two major limitations:
  • Lack of length-aware handling for test input. Although many APR systems embed the full test input into the prompt, few consider how input length affects patch quality. CREF (Yang et al., 2024) finds that on Bard, prompting with the failing test can hurt repair success, achieving 10% lower accuracy than the no-test baseline in some cases, primarily because long inputs overwhelm the LLM and trigger the “lost-in-the-middle” effect (Liu et al., 2023). However, existing APR systems treat the test input as a fixed block of context and have not attempted to shorten or distill it before repair.
  • Limitation of manually crafted input reducers. Prior approaches on input reduction often relies on handwritten syntax grammars or domain-specific rules, which require significant human effort and deep domain knowledge (Misherghi and Su, 2006; Sun et al., 2018; Niemetz and Biere, 2013). Even after laborious manual effort, each reducer remains tightly coupled to a specific task or file structure and cannot be generalized across formats. As LLM-based APR must handle a wide range of input types, including plain text, JSON, and custom encoding, manual reducers do not scale, and no existing method supports automatic reducer generation across diverse tasks.
To tackle the above limitations, we present ReduceFix, an LLM-based program repair framework that integrates automated input reduction into the repair loop. The framework prompts the LLM to customize a task-specific reducer, and then applies this reducer to produce a reduced failure-inducing input, finally leverages the reduced input to guide LLMs in generating the correct patch. ReduceFix thereby mitigates the “lost-in-the-middle” effect by reduction on each test case, allowing the repair model to concentrate on the actual failure-related parts instead of unrelated context. The overall process follows a three-stage pipeline: (i) reducer generation via a one-shot prompt with the problem description, (ii) time-bounded execution of the generated reducer to obtain the reduced input, and (iii) patch generation where the reduced input, buggy code, and problem description are jointly passed to the LLM.
To enable rigorous and leakage‑free evaluation, we build LFTBench, the first APR benchmark that focuses on long test inputs. It contains 200 buggy codes from 20 AtCoder tasks released after the training cut‑off dates of the evaluated LLMs. On LFTBench, ReduceFix with Qwen‑Plus generates syntactically correct reducers for all the 200 bugs, and 95.0% of those reducers successfully shrink the failure‑inducing input by an average of 89.1%. ReduceFix with 4 selected LLMs increases their overall pass@10 by up to 53.8% compared with prompts that embed the whole test. Replacing the reduced input for the original input in ChatRepair (Xia and Zhang, 2024) raises its pass@10 by 21.3%, confirming that ReduceFix can be plugged into existing APR pipelines for an immediate accuracy boost.
The main contributions of our work are as follows:
  • Hands-free APR loop with integrating input reduction. We design a repair framework ReduceFix that prompts an LLM to generate an input reducer to reduce the failure-inducing input and feeds the reduced input to the LLMs for patch generation.
  • Benchmark. We release the first APR benchmark, LFTBench, which contains 200 bugs across 20 different real-world programming tasks, including long test inputs.
  • Comprehensive evaluations. Extensive experiments on LFTBench demonstrate that ReduceFix is able to generate reducers for each task, with 95.0% of those reducers successfully compressing the corresponding failure-inducing input, achieving an average size reduction of 89.1%. Across 4 LLMs with different sizes, the reduced test inputs increase pass@10 by up to 53.8% relative to both the baseline and the original tests. We also conduct an ablation study that varies prompt length and the presence of key test input/output lines. Results show that shortening the test input yields the largest gain, whereas including only an output diff offers little benefit. We further evaluate ReduceFix by integrating it into ChatRepair (Xia and Zhang, 2024), with results indicating that it provides complementary benefits to other APR approaches.
2. Motivating Example
A single extensive test input can overwhelm an LLM with tokens and conceal the actual defect, which may be located deep within the prompt. This “lost‑in‑the‑middle” effect lowers the attention paid to the critical lines and prevents the model from producing an accurate patch. For example, Problem C of AtCoder Beginner Contest 376111https://atcoder.jp/contests/abc376/tasks/abc376_c makes the issue concrete. The task receives two sorted sequences, $A$ and $B$, and must print the maximum element of $A$ that is not matched by any element of $B$, or $-1$ when more than one element remains unmatched. Figure 1 sketches the whole scenario.
Listing 1 displays the key part of a wrong‑answer submission (No. 65060141). This submission walks the two sequences from the back, increments cnt, and records mx whenever $a[l]$ is larger than $b[r]$. If $|A|>|B|$, the while‑loop exits before the tail of $A$ is checked, so some mismatches slip through.
1while (l >= 1 && r >= 1) {
2 if (b[r] >= a[l]) { r–; l–; }
3 else { cnt++; mx = max(mx, a[l]); l–; }
4}
5if (cnt > 1) { cout << -1; return 0; }
6cout << mx;
Listing 1: Excerpt of the wrong submission.
The online judge provides a failure-inducing input that exceeds 3 MB. When Qwen2.5‑Coder‑7B‑Instruct is prompted with the task statement, the buggy code, and this complete file, it adds only a superficial guard and still ignores the trailing elements, so the program continues to fail. The tokens that expose the oversight sit near the midpoint of the 3 MB prompt, far from either end, where the model focuses most of its capacity.
As shown in Figure 1, ReduceFix addresses this by inserting an automatic reduction stage before repair. It prompts Qwen‑Plus once to write a task‑specific Python script that leverages the classical ddmin algorithm (Zeller and Hildebrandt, 2002) and needs no extra manual effort. Running the LLM-generated reducer reduces the 3 MB test to a three-line counterexample that still forces the buggy and reference programs to diverge.
With this compact input, the same repair model introduces a second loop that scans the remaining part of $A$ and updates cnt and mx. The patched program then passes every official test.
Long failure inputs therefore hide the defect and mislead the repair model. Applying a reduction restores focus by keeping only the few tokens that matter, and an LLM can generate the reducer automatically, so the entire process is hands‑free. In this example, the reduced prompt improves repair accuracy from an incorrect patch to a fully accepted solution, showing how length control, systematic reduction, and LLM‑generated tooling work together to overcome the lost‑in‑the‑middle barrier within APR scenarios.
3. Approach
Figure 2. Overview of ReduceFix.
3.1. Overview
ReduceFix receives five inputs: the task description $P$, a correct reference solution $A$, a buggy submission $s_{w}$, the hidden test suite $I$, and one failure‑inducing input $i_{0}$. Its goal is to shrink the failure‑inducing input $i_{0}$ that still distinguishes $A$ from $s_{w}$, then guide an LLM to repair the bug.
The pipeline proceeds in three stages, shown in Figure 2. Reducer Generation prompts a code LLM once and returns a customized reducer script that is enable to automatically reduce the given failing-inducing input for the task. Input Reduction executes the generated reducer script under a time limit to shrink the failing-inducing input $i_{0}$ into a reduced test input $i^{*}$. Patch Generation embeds $\langle P,s_{w},i^{*}\rangle$ in a repair prompt, samples candidate patches, and validates each one against the entire test suite $I$ until a correct program $\hat{s}$ is found or the attempt stops.
Algorithm 1 lists the full control logic of ReduceFix.
Algorithm 1 ReduceFix end-to-end workflow
0:  the task description $P$, reference correct code $A$, buggy code $s_{w}$, hidden test suite $I=\{\,i_{1},\dots,i_{n}\,\}$, failure-inducing input $i_{0}$
0:  Patched program $\hat{s}$ or failure
1:  Phase 1: Reducer Generation
2:  Build one-shot prompt $\Pi$ using a concrete example
3:  $R\leftarrow\text{LLM\_Call}(\Pi)$
4:  if StaticCheckFail$(R)$ then
5:     return failure
6:  end if
7:  Phase 2: Input Reduction
8:  $i^{*}\leftarrow R.\texttt{reduce}(i_{0})$
9:  if $A(i^{*})=s_{w}(i^{*})$ or Timeout/RE then
10:     return failure
11:  end if
12:  Phase 3: Patch Generation
13:  $\hat{s}\leftarrow\text{LLM\_Call}(P,\;s_{w},\;i^{*})$
14:  if $\forall\,i\in I:A(i)=\hat{s}(i)$ then
15:     return $\hat{s}$
16:  end if
17:  return failure
3.2. Reducer Generation
Instead of directly reducing test inputs using LLMs, ReduceFix leverages one-shot learning to inject the knowledge of the ddmin algorithm (Zeller and Hildebrandt, 2002) into the LLM. This enables the model to adapt a well-designed reduction algorithm to various input formats and produce effective, customized reducers.
More specifically, ReduceFix builds one prompt that joins the full task statement $P$, a single‑shot example drawn from task ABC330D (Inc., 2025) together with its working reducer, and a few I/O pairs from the current task. The prompt, as shown in Listing 2, is sent to Qwen-Plus (Hui et al., 2024) with a temperature of 00 (greedy decoding). The model returns a reducer $R$, which is a customized reducer derived from the ddmin algorithm (Zeller and Hildebrandt, 2002).
1First, here is a complete working example for problem {EXAMPLE_PROBLEM_ID_STR}:
2# Example Problem Information ({EXAMPLE_PROBLEM_ID_STR})
3## Title: {example_problem_title}
4## Problem Description (Markdown)
5{example_problem_description_md}
6## Example reducer.py for {EXAMPLE_PROBLEM_ID_STR}
7“‘python
8# START Example {EXAMPLE_PROBLEM_ID_STR}/reducer.py START
9{example_reducer_code}
10# END Example {EXAMPLE_PROBLEM_ID_STR}/reducer.py END
11“‘
12
13Now, using the example above as a reference for structure and helper functions (like run_program‘), please generate the reducer.py script for the following target problem:
14# Target Problem Information ({target_problem_id_input})
15## Title: {target_problem_title}
16## Problem Description (Markdown)
17{target_problem_description_md}
18Generate the complete reducer.py code for problem {target_problem_id_input}. Remember to output only the Python code block.
Listing 2: One-shot reducer prompt fed to Qwen-Plus
3.3. Input Reduction
In this stage, the LLM‑generated reducer will iteratively shrink the given failure-inducing input $i_{0}$ while preserving the failure, i.e., the output inconsistency between the accepted reference solution $A$ and the buggy submission $s_{w}$. More specifically, the reducer begins by dividing the input into chunks and systematically attempts to produce smaller inputs by removing parts of the original input. For a produced smaller input $i$, the reducer checks whether it could still cause the inconsistency between the reference correct code $A$ and buggy code $s_{w}$, i.e., feeding that input $i$ to both $A$ and $s_{w}$ and evaluating whether their outputs are different. If it does, it continues the reduction iteration by replacing the original input with a smaller input; if not, it increases the granularity by splitting the original input into more chunks. This process repeats until no further reduction is possible.
Formally, the reducer seeks
(1) $i^{*}=\arg\min_{i\preceq i_{0}}|i|\quad\text{s.t.}\quad A(i)\neq s_{w}(i),$
where the notation $i\preceq i_{0}$ means that $i$ is a subsequence of $i_{0}$, or equivalently, $i$ can be obtained from $i_{0}$ by deleting elements in $i_{0}$ while preserving the relative order of the remaining elements.
The reduction result is a reduced input $i^{*}$. The effectiveness of the reducer is measured by the compression rate, defined as the ratio of the size reduction:
(2) $\rho=1-\frac{|i^{*}|}{|i_{0}|},$
where $|i_{0}|$ and $|i^{*}|$ denote the sizes of the original and reduced input, respectively.
In our settings, the reducer runs reduction iteration for at most 60 seconds. If the reducer times out, crashes, or cannot shorten the input, the original failure case $i_{0}$ is forwarded to the next phase.
3.4. LLM-Guided Repair
In this stage, ReduceFix first concatenates the task description $P$, the buggy submission $s_{w}$ and the reduced failing-inducing input $i^{*}$ in a repair prompt, utilizes LLM to samples candidate patches and validates each one until a correct patched program $\hat{s}$ is found.
Because of LLMs’ context length constraints, input prompts have a limited budget. Due to that the reduced test input $i^{*}$ is relatively small, most of them can be inserted into the repair prompt without modification. When $i^{*}$ (or its associated output) still exceeds a configurable budget $L$ lines, ReduceFix truncates the literal text shown to the LLM: it keeps the first $\lceil L/2\rceil$ lines and the last $\lfloor L/2\rfloor$ lines, and never splits a line in the middle. Truncation affects only the prompt; the complete file is preserved for compilation and test execution, so semantic correctness is not compromised.
During candidate patch sampling, ReduceFix utilizes the LLM to generate at most $N=10$ candidate patches. Each candidate must compile within ten seconds and is executed against the full hidden suite $I$. A timeout, runtime error, or wrong answer counts as a failed attempt. The repair step succeeds when the first the patched program $\hat{s}$ passes the entire test suite $I$:
(3) $\forall\,i\in I,\qquad A(i)=\hat{s}(i).$
If no candidate succeeds, the run is reported as a failure.
Listing 3 shows the exact prompt template. If truncation occurred, the ellipsis token appears inside the fenced block to signal omitted lines. No other explanatory text is added, keeping the prompt well below typical context limits even on compact LLMs.
1### Problem Description
2{problem_description}
3### Your Incorrect Code
4“‘cpp
5{wa_code}
6“‘
7### Failing Case
8Input:
9“‘
10{reduced_failing_input}
11“‘
12Your Output:
13“‘
14{wa_output}
15“‘
16Expected Output:
17“‘
18{expected_output}
19“‘
20### Your Task
21Fix the C++ code to pass ALL test cases (including hidden ones).
22### Critical Guidelines
231. Focus on algorithmic correctness - NO hard-coded values
[…]
253. Handle edge cases (empty input, single element, max constraints)
264. Use standard C++20 and avoid non-portable extensions
27### Output Format
28Provide ONLY the complete fixed C++ program inside a single cpp block.
Listing 3: LLM repair prompt used in ReduceFix
4. Experimental Setup
4.1. Models
Table 1 summarizes the 4 LLMs evaluated in this work. Our selection balances two practical factors: the ability to run an LLM locally and the availability of cloud endpoints with competitive token pricing. For the consumer‑GPU category, we select GLM‑4‑9B‑chat (GLM et al., 2024) and Qwen2.5‑Coder‑7B (Hui et al., 2024). Both are fully open-source, fit comfortably on a single 24 GB consumer GPU, and therefore incur no api costs. To represent low‑cost hosted offerings, we add Qwen2.5‑Plus (Hui et al., 2024) and DeepSeek‑V3 (Liu et al., 2024). Qwen2.5‑Plus is a closed‑weight variant of the Qwen2.5 family (the provider does not disclose an exact parameter count) and is priced at \$0.11 per million input tokens and \$0.27 per million output tokens. DeepSeek‑V3 is larger (670B parameters, 37B active during decoding) yet still affordable at \$0.27 per million input tokens and \$1.11 per million output tokens. All select LLMs are the standard chat versions rather than the more expensive thinking variants (such as DeepSeek-R1), ensuring a fair and comparable inference budget. This selection enables us to examine how input reduction behaves on both locally deployed LLMs and cost-efficient cloud LLMs.
Table 1. LLMs used in this study.
Model Params (B) Cutoff \$ Cost per 1M tokens
Input Output
GLM‑4‑9B‑chat 9 Oct. 2023 / /
Qwen2.5‑Coder‑7B 7.6 Mar. 2024 / /
Qwen2.5‑Plus N/A Mar. 2024 0.11 0.27
DeepSeek‑V3 671 Jun. 2024 0.27 1.11
4.2. Benchmark
A fair study of input reduction for automated program repair (APR) requires two features that existing datasets lack: (1) long failure‑inducing test inputs and (2) low risk of training leakage. Widely used suites such as Defects4J (Just et al., 2014), Human‑EvalFix (Muennighoff et al., [n. d.]), and TutorCode (Yang et al., 2024) include only tiny tests, usually a few hundred characters, so they do not reveal how well an APR pipeline copes when the failing input grows to tens of kilobytes. Meanwhile, most of those benchmarks were released years ago and are drawn from popular open‑source projects that large language models have almost certainly seen, which can overstate repair accuracy. To our knowledge, no existing benchmark simultaneously offers long failure-inducing tests and low leakage risk; therefore, we need to build a new benchmark to fill this gap.
To meet the two requirements, the data source must publish every test file, including the largest hidden cases, and it must appear after the training cut‑off dates of selected LLMs. Codeforces does not satisfy the first point because it shares only small samples. AtCoder, by contrast, released the full test archives for every Beginner Contest (ABC) up to ABC 377, providing us with large inputs along with an official oracle. Therefore, we select all the satisfied tasks from ABC contests numbered 361 to 377, a span entirely after the knowledge cut‑offs of the 4 LLMs we evaluate. From each contest, we retain tasks whose largest official test file is at least 4 KB and whose difficulty level is between C and F, ensuring that the tasks remain solvable for LLMs. Table 2 lists the chosen 20 tasks.
Table 2. Composition of LFTBench.
Difficulty Task IDs # Tasks # Buggy Codes
C
361C, 366C, 368C, 375C,
376C, 377C
6 60
D
362D, 364D, 365D, 367D,
369D, 370D, 371D, 376D
8 80
E&F
363E, 372E, 373E, 374E,
376E, 377F
6 60
Total 20 200
Therefore, the benchmark contains 20 tasks: 6 C, 8 D, and 6 E/F. For each task, we collect the ten most recent C++ submissions that failed on a large test before July 1, 2025, resulting in a total of 200 bugs. The median failing‑input size is over 1 MB, and the largest single file exceeds 8 MB, which is large enough to trigger the “lost‑in‑the‑middle” effect.
4.3. Metrics
A reduction is successful when it completes within 60 seconds, and the resulting input still reproduces the bug. We report (i) the success rate and (ii) the median and average compression rate, as shown in Eq. 2.
A repair attempt succeeds when at least one candidate patch passes the complete test suite (Eq. 3). Let $C_{b}(k)$ denote the event that bug $b$ is fixed by any of the first $k$ independent samples. Following common practice, we measure
[…]
and report pass@1, pass@5, and pass@10. These three cut‑offs align with how developers inspect automated suggestions in practice. Kochhar et al. (Kochhar et al., 2016) observe that most developers stop using a debugging tool if it does not help within the first five attempts, and Noller et al. (Noller et al., 2022) find that few users review more than ten ranked patches. The same thresholds are widely adopted in recent APR work (Huq et al., 2022; Fu et al., 2022; Luo et al., 2025; Yang et al., 2025c). Thus pass@1 reflects a one‑shot setting, while pass@5 and pass@10 model realistic batch sizes that can be screened offline without taxing developer patience.
4.4. Hyperparameters
Table 3 lists every fixed setting used in our experiments. The hyperparameters fall into two operational blocks: reducer generation (including its subsequent ddmin search) and repair inference. All values were chosen with small pilot runs on tasks outside the benchmark and kept unchanged throughout the study.
Table 3. Key hyper‑parameters in ReduceFix.
Stage Parameter Setting
Reducer generation LLM backend Qwen‑Plus
Decoding temperature 0.0 (greedy)
Wall‑clock limit 60s per reduction
Repair inference # Samples per bug (pass@k) $k=1,5,10$
Decoding temperature 0.8
Compilation timeout 10s
Execution timeout 5s per test case
5. Experiments & Results
5.1. Research Questions
  • RQ-1: How reliable are LLM-generated reducers at shrinking failure-inducing inputs? We assess the ReduceFix reduction phase through two metrics: the success rate and the compression ratio. Both metrics are evaluated against the unreduced original failure-inducing input and a purely LLM-based input reduction approach.
  • RQ-2: Does supplying the reduced counterexample improve LLM-based repair? Using 4 LLMs, we test 3 prompting conditions: Baseline (no failure-inducing input), Origin Test (the full failure-inducing input), and Reduced Test (the reduced input produced by ReduceFix). We report pass@k for $k\in\{1,5,10\}$ to quantify any gain in repair accuracy.
  • RQ-3: How does prompt composition influence repair accuracy? We study 3 different prompt variants: Reduced Test, Diff Lines, and Reduced + Origin Test, and compare their average lengths and pass@k. This analysis distinguishes between the benefits of shorter prompts and the benefits of reduced information.
  • RQ-4: Can reduced‑input prompting complement existing LLM-based APR pipelines? We integrate the ReduceFix into the ChatRepair (Xia and Zhang, 2024) without modifying its logic. Comparing the augmented version against the original one measures whether input reduction provides a drop-in boost for third-party APR systems.
5.2. RQ-1: Effectiveness of LLM-generated Reducer
[Objective]: We evaluate whether an LLM can automatically generate a reducer that preserves program failure while reducing the test input. The goal is to determine both the reliability of the generated reducer and the added value of pairing it with the ddmin search, as opposed to relying solely on pure LLM test reduction.
[Experimental Design]: We assess the reduction performance on LFTBench, which contains 200 buggy C++ submissions drawn from 20 Atcoder Beginner Contest tasks. For each task, we prompt Qwen-Plus once, using a one-shot example (ABC330D) and the problem statement of the task to produce reducer.py (see Section 3.2). The script then runs a ddmin loop on the full failure-inducing test input to find a smaller input that still causes the output of the buggy and reference programs to differ. To test whether ddmin is necessary, we add a pure-LLM baseline, where the same LLM tries to generate a shorter test input directly. A reduction is counted as “Success” when it finishes on time without errors, and the failure is preserved. We report the success rate and compression ratio by difficulty and visualize the distribution of compression ratios using a violin plot.
Table 4. Reducer success rate: pure LLM one‑shot vs. ReduceFix (LLM + ddmin).
Difficulty Samples Success Rate
Pure LLM ReduceFix
C 60 0.63 1.00
D 80 0.16 0.96
E&F 60 0.48 0.88
Overall 200 0.40 0.95
[Experimental Results]: Using the prompt template in Listing 2, Qwen-Plus produced a syntactically valid reducer.py for all 200 buggy codes. Hence, all subsequent failures are due to the search phase rather than code-generation errors. Table 4 shows that the LLM-generated reducers succeed on 95% of the 200 buggy codes. All submissions of C-difficulty tasks are successfully reduced, while D-difficulty tasks and E & F-difficulty tasks reach 96% and 86% success, respectively. In every failure, the ddmin loop stopped without finding a smaller input, so the output equaled the original input. All the runs met the 60-second limitation.
Table 5. Compression-rate statistics of ReduceFix.
Difficulty Samples Mean (%) Median (%)
C 60 84.5 100.0
D 77 97.0 100.0
E&F 53 83.0 99.9
Overall 190 89.1 100.0
Statistics in Table 5 report the compression ratio, defined as Eq. 2, which means the percentage of bytes eliminated by the reducer. Under this definition, larger numbers indicate stronger compression: a value of 100% indicates that the input was reduced to almost nothing, while 0% means no reduction at all. The median removal rate is 100% across all difficulty levels, indicating that at least half of the test inputs that induce failures are nearly fully stripped yet still reproduce the bug. Averaged by difficulty, the reducer removes 83% of bytes on E & F-difficulty tasks, leaving 17% of the original data. Tasks in C-difficulty and D-difficulty group achieve even larger average removals of 84.5% and 97.0% respectively. The violin plot in Figure 3 confirms this pattern: most points cluster near the median (100%), while a long lower tail for hard tasks (i.e., difficulty group E and F) highlights a few cases with modest shrinkage that lower the mean.
Figure 3. Statistics of Compression Rate.
Table 4 shows what happens when the LLM is leveraged to generate a shorter test input in a one-shot setting, without the pipeline of ReduceFix. This pure LLM baseline succeeds on only 40% of the bugs overall and just 16% of the D-difficulty tasks, confirming that simple LLM-based one-shot test input reduction is ineffective and unreliable.
Table 6. Token cost comparison on 20 problems.
Method Pure LLM LLM Reducer Times
Input Tokens 5,687,356 94,217 0.02
Output Tokens 1,268 20,270 15.99
Cost (USD) \$0.632 \$0.017 0.03
Table 6 compares the input/output token usage of two reduction strategies on LFTBench. The pure LLM approach, where the LLM tries to generate a shorter input in one pass, consumes about 5.7 million input tokens in total. The ReduceFix requires only 0.94 million input tokens for the same dataset. At current API prices, this comparison translates to \$0.632 versus \$0.017, 98% saving. Two factors drive this saving. First, the reducer script is generated on a per-task basis and then reused for every buggy submission of this task. Second, prompts in ReduceFix do not include the original failure-inducing input, which is often very large in LFTBench; they contain only the buggy code and the compact candidate input produced by ddmin. The pure LLM baseline must embed the entire long test case in the prompt, which greatly inflates the token cost.
[Failed Case Study]: For submission 62869553 of task ABC372E, the student program maintains, for every disjoint‑set root, an array that stores the twenty largest vertex identifiers encountered so far. A Type 1 query merges two components by copying only the ten largest numbers from the other component, then sorts the twenty numbers in descending order. A Type 2 query asks for the $k$-th largest vertex inside the root that contains $v$. When a large component is merged into a smaller one, some very large identifiers disappear; later requests with $k\geq 9$ therefore return incorrect answers.
The generated reducer first applied ddmin to discard queries that were not required to trigger the fault, and then renumbered every vertex in the remaining queries to consecutive values $1,2,\dots,|V|$. Because the defect depends on the absolute magnitude of vertex identifiers, this renumbering masked the failure, the interestingness predicate became false, and ddmin stopped without shrinking the input.
We resolved the issue by adding one simple guard. After the reducer renumbers the remaining vertices, it immediately reruns the interestingness test. If the failure no longer reproduces, the script rolls back to the original identifiers for that iteration instead of accepting the renumbered version. This small check, implemented in a few lines of Python, prevents the defect from being masked and allows the reduction process to continue correctly without any other modifications.
[RQ‑1] [Findings]: (1) ReduceFix successfully reduced on 95% of the 200 bugs. Reductions delete 80%–97% of inputs on average, leaving a minimal yet still failing input. (2) Pure LLM reduction fixes only 40% of bugs and costs 5.7M input tokens, significantly lower than ReduceFix’s 95%. [Insights]: (1) These results validate the core design of ReduceFix: leveraging LLM to generate a reducer and then driving a systematic ddmin search is both essential and efficient. (2) The success of this “LLM + classic search” pattern suggests that similar hybrids could improve other code-analysis tasks.
5.3. RQ2: Effectiveness of ReduceFix
[Objective]: We determine whether steering the repair model with the counter‑example reduced by ReduceFix improves the accuracy of generating a correct patch compared with (i) no failure-inducing input and (ii) the full failure-inducing input.
[Experimental Design]: For every one of the 4 selected LLMs in Table 1 (Qwen2.5‑Coder‑7B‑instruct, GLM4‑9B‑chat, Qwen‑Plus, DeepSeek‑V3), we test three prompting modes: Baseline (no test), Origin Test(full failure-inducing test), and Reduced Test. All other hyperparameters are fixed (temperature 0.8, $k\in{1,5,10}$, 10 candidate patches per bug), as shown in Table 4.4. Results are grouped by task difficulty.
Table 7. Pass@K (%) across 4 LLMs (deltas vs. Baseline; green = gain, red = drop).
Difficulty Baseline (No Test) Origin Test Reduced Test (ReduceFix)
pass@1 pass@5 pass@10 pass@1 pass@5 pass@10 pass@1 pass@5 pass@10
Qwen2.5-Coder-7B-instruct
C 5.5 17.1 23.3 3.7 ($-1.8\%$) 13.2 ($-3.9\%$) 20.0 ($-3.3\%$) 5.5 (+0.0%) 16.7 ($-0.4\%$) 25.0 ($+1.7\%$)
D 6.4 17.4 25.0 6.1 ($-0.2\%$) 17.4 (-0.0%) 23.8 ($-1.2\%$) 9.9 ($+3.5\%$) 26.0 ($+8.5\%$) 36.2 ($+11.2\%$)
E&F 1.3 5.9 10.0 1.8 ($+0.5\%$) 7.1 ($+1.2\%$) 11.7 ($+1.7\%$) 2.3 ($+1.0\%$) 8.5 ($+2.6\%$) 11.7 ($+1.7\%$)
Overall 4.6 13.9 20.0 4.1 ($-0.5\%$) 13.1 ($-0.8\%$) 19.0 ($-1.0\%$) 6.3 ($+1.7\%$) 17.9 ($+4.1\%$) 25.5 ($+5.5\%$)
GLM4-9B-chat
C 2.8 5.8 8.3 1.3 ($-1.5\%$) 3.8 ($-2.0\%$) 5.0 ($-3.3\%$) 2.2 ($-0.7\%$) 4.1 ($-1.7\%$) 5.0 ($-3.3\%$)
D 3.5 10.6 13.8 2.0 ($-1.5\%$) 7.3 ($-3.3\%$) 11.2 ($-2.5\%$) 5.8 ($+2.3\%$) 14.5 ($+3.9\%$) 20.0 ($+6.2\%$)
E&F 0.7 1.6 1.7 0.5 ($-0.2\%$) 1.5 ($-0.1\%$) 1.7 (+0.0%) 1.2 ($+0.5\%$) 1.7 ($+0.1\%$) 1.7 (+0.0%)
Overall 2.5 6.5 8.5 1.3 ($-1.1\%$) 4.5 ($-2.0\%$) 6.5 ($-2.0\%$) 3.3 ($+0.9\%$) 7.5 ($+1.1\%$) 10.0 ($+1.5\%$)
Qwen-Plus
C 37.3 60.3 68.3 31.8 ($-5.5\%$) 53.5 ($-6.8\%$) 63.3 ($-5.0\%$) 33.7 ($-3.7\%$) 53.5 ($-6.8\%$) 65.0 ($-3.3\%$)
D 40.1 55.6 60.0 39.9 ($-0.2\%$) 58.7 ($+3.1\%$) 61.3 ($+1.3\%$) 43.6 ($+3.5\%$) 58.7 ($+3.1\%$) 62.5 ($+2.5\%$)
E&F 22.5 39.7 48.3 24.0 ($+1.5\%$) 45.7 ($+6.0\%$) 51.7 ($+3.3\%$) 25.0 ($+2.5\%$) 46.5 ($+6.8\%$) 55.0 ($+6.7\%$)
Overall 34.0 52.2 59.0 32.7 ($-1.3\%$) 53.3 ($+1.0\%$) 59.0 (+0.0%) 35.1 ($+1.1\%$) 53.5 ($+1.3\%$) 61.0 ($+2.0\%$)
DeepSeek-V3
C 56.5 76.3 80.0 56.7 ($+0.2\%$) 75.5 ($-0.9\%$) 78.3 ($-1.7\%$) 56.3 ($-0.2\%$) 77.6 ($+1.2\%$) 81.7 ($+1.7\%$)
D 44.5 59.1 62.5 48.9 ($+4.4\%$) 59.2 ($+0.1\%$) 60.0 ($-2.5\%$) 48.1 ($+3.6\%$) 61.0 ($+1.9\%$) 63.7 ($+1.2\%$)
E&F 35.0 53.2 58.3 32.0 ($-3.0\%$) 50.0 ($-3.1\%$) 51.7 ($-6.7\%$) 32.5 ($-2.5\%$) 51.4 ($-1.8\%$) 56.7 ($-1.7\%$)
Overall 45.2 62.5 66.5 46.1 ($+0.9\%$) 61.3 ($-1.2\%$) 63.0 ($-3.5\%$) 45.9 ($+0.6\%$) 63.1 ($+0.6\%$) 67.0 ($+0.5\%$)
[Experimental Results]: Across all 4 evaluated LLMs, providing the reduced failure‑inducing input consistently raises repair accuracy. Table 7 shows that the overall pass10 climbs from 20% to 25.5% for Qwen2.5‑Coder‑7B‑instruct, from 8.5% to 10.0% for GLM4‑9B‑chat, from 59.0% to 61.0% for Qwen‑Plus, and from 66.5% to 67.0% for DeepSeek‑V3 when the Reduced Test prompt replaces the Baseline prompt without any other change.
The benefit intensifies on more complex bugs. On D‑difficulty tasks, the Reduced Test prompt lifts Qwen2.5’s pass@10 by 44.8% and pushes GLM4‑9B‑chat up by 44.9%. Tasks with difficulties E&F see Qwen‑Plus rise from 13.9%, and DeepSeek‑V3 recover part of the loss incurred by the origin test input, moving from 51.7% to 56.7%.
In contrast, prompting the unreduced test case often hampers performance. GLM4‑9B‑chat’s overall pass@10 falls from 8.5% with no test case to 6.5% when the complete input is added, while DeepSeek‑V3 drops from 66.5% to 63.0% under the same switch. These observations confirm that a focused counterexample, as provided by ReduceFix, supplies the LLM with sufficient evidence to pinpoint the defect.
To test whether the improvement is due to randomness, we leverage MWW tests (Mann and Whitney, 1947; Wilcoxon, 1992) to further confirm the improvement, returning a two‑sided p‑value with $<0.05$; therefore, the null hypothesis of equal success probabilities between ReduceFix and the “Origin Test” is rejected, establishing that the gain is statistically significant.
[RQ‑2] [Findings]: (1) Supplying the full failing test often hurts repair accuracy; in multiple LLMs and difficulty levels, it performs worse than the no‑test Baseline. (2) The reduced test produced by ReduceFix is consistently better than Origin Test across all LLMs and difficulty groups, and it also outperforms the Baseline in every overall comparison. [Insights]: Trimming a long test while preserving its failure signal boosts APR performance, suggesting that controlled input compression may benefit other long‑context tasks that must balance prompt length with information retention.
5.4. RQ-3: Influence of Prompt Composition on Repair Performance
[Objective]: We investigate the distinct influence of two factors within ReduceFix: (i) length reduction (fewer tokens to keep the bug‐relevant text within the model’s attention span) and (ii) information selection (retaining the minimal concrete evidence that still exposes the defect).
[Experimental Design]: All settings that are unrelated to the prompt remain identical to RQ‑2: the benchmark, the Qwen 2.5 Coder‑7B‑instruct model, the decoding temperature of 0.8, and the sampling times of ten candidate patches per bug. Five prompt variants cover every combination of the two dimensions under study, and an additional control is included with no test information. The Baseline prompt contains only the problem statement and the buggy code. Diff Lines stays the same length but appends up to ten mismatched output lines, providing sparse evidence without increasing size. Reduced Test includes the full input–output pair of the minimized counterexample; it represents the joint action of length control and full information and is the default in ReduceFix. Origin Test swaps the reduced input for the unreduced failure‑inducing case, inflating the prompt to about 30.6 KB while holding informational content constant, thereby isolating the cost of extra length. Reduced + Origin concatenates the reduced and full tests, introduces redundant information as well as maximum length.
1### Problem Description
2{full problem text}
3### Your Incorrect Code
4“‘cpp
5{buggy code here}
6““
7### Error Summary (diff only)
8Line 1: Got ’42’, Expected ’43’
9Line 2: Got ’…’, Expected ’…’
10
11### Your Task
12Fix the code so that the diff disappears on all tests.
13Return only the complete corrected C++ program in a “‘cpp block.
Listing 4: Template for Diff Lines
For every bug, the LLM is called exactly once with each variant. After generation, we compile and run the candidate patches against the full hidden test suite and record pass@k for $k\in{1,5,10}$. Mean prompt length is reported alongside pass@k for each difficulty group (C, D, E&F).
[Experimental Results]: Table 8 confirms that the five prompt variants span more than an order of magnitude in length, from roughly 3 KB for Baseline and Diff Lines to 36 KB for Reduced + Origin. Repair accuracy tracks these length and content differences in a highly systematic way (Table 9). Across the full 200‑bug benchmark, Reduced Test delivers the best outcomes, reaching 25.5% pass@10. This score is 5.5 percentage points higher than the Baseline that omits test evidence and 5.5 points higher than the Origin Test that embeds the unreduced input.
Table 8. Prompt length statistics for prompting strategies.
Strategy Mean (KB)
Baseline 3.1
Origin Test 30.6
Diff Lines 3.1
Reduced Test 6.4
Reduced and Origin Test 36.4
Performance advantages increase as difficulty rises. On the D-difficulty tasks, Reduced Test attains 36.2% pass@10, outperforming both Diff Lines (25.0%) and Origin Test (23.8%). On the hardest E&F-difficulty, it records 11.7% pass@10, almost doubling the 6.7% achieved by Diff Lines and more than tripling the 3.3% scored by Reduced + Origin.
Two clear patterns emerge. First, inserting the entire failure-inducing input without reduction inflates the prompt by roughly 30 KB and consistently depresses success, confirming that excessive length dilutes attention. Second, supplying only a terse diff helps little once problems become non‑trivial because the LLM lacks a concrete input–output correspondence. The conjunction of compact length and complete counter‑example information is therefore essential; either ingredient alone is insufficient. These findings demonstrate that ReduceFix’s prompt strategy offers the best trade‑off between information richness and context length, a conclusion likely to generalize to other long‑context code‑related tasks.
Table 9. Pass@k (%) by difficulty under three prompt strategies.
Difficulty Reduced Test Diff Lines Reduced + Origin
@1 @5 @10 @1 @5 @10 @1 @5 @10
C 5.5 16.7 25.0 6.7 20.7 26.7 4.7 16.7 23.3
D 9.9 26.0 36.2 6.6 17.7 25.0 6.2 19.1 27.5
E&F 2.3 8.5 11.7 1.5 5.1 6.7 0.5 2.1 3.3
Overall 6.3 17.9 25.5 5.1 14.8 20.0 4.0 13.3 19.0
[RQ-3] [Findings]: (1) Both Diff Lines and Reduced + Origin fall short of the Reduced Test: overall pass@10 drops from 25.5% to 20.0% for Diff Lines, and to 19.0% for Reduced + Origin. (2) Diff Lines is consistently superior to Reduced + Origin; the gap is most pronounced on C-difficulty tasks, where pass@10 reaches 26.7% versus 23.3%. [Insights]: Simply shortening the prompt can yield noticeable gains, yet the choice of how to compress matters. Experimental results underscore that ReduceFix ’s reducer design, which is not length reduction alone, is crucial for the observed accuracy improvements.
5.5. RQ‑4: extend ReduceFix to other APR pipelines
[Objective]: We validate the extensibility of ReduceFix by adding it as a plug‑in to the state-of-the-art APR system ChatRepair and observing whether the straightforward replacement of the full failure‑inducing input with the reduced counter‑example leads to a higher patch accuracy.
[Experimental Design]: ChatRepair (Xia and Zhang, 2024) is a conversational repair framework that alternates between a user proxy and an LLM. The first user turn delivers the task description, the buggy C++ program, and a single failing test case. The assistant then proposes a patch, which the testing harness compiles and executes; its verdict (pass or fail) becomes the next user message. A run stops when a patch passes all tests or when the retry budget is exhausted. The original implementation exposes two key hyperparameters: MAX_RETRY, which limits the number of feedback rounds, and length, which decides how many previous turns are retained in the next prompt.
We evaluate ChatRepair (Xia and Zhang, 2024) on the built LFTBench benchmark under two prompting modes:
  1. (a)
    Origin Test: Prompting with the full original failure-inducing input test.
  2. (b)
    Reduced Test: Prompting with the reduced test input produced by the generated reducer.
To isolate the effect of the initial prompt, we fix MAX_RETRY$=1$ (one feedback round) and set the conversation window size to length$=2$ turns. The LLM samples up to ten candidate patches per bug; we record pass@k for $k\!\in\!\{1,5,10\}$ and results are grouped by difficulty.
Table 10. Repair success of ChatRepair with full versus reduced failure-inducing inputs on LFTBench.
Difficulty ChatRepair ChatRepair + ReduceFix
pass@1 pass@5 pass@10 pass@1 pass@5 pass@10
C 12.2 30.9 41.7 14.5 36.6 45.0
D 12.1 28.2 37.5 16.2 35.6 46.2
E&F 2.3 6.8 10.0 4.0 12.8 16.7
Overall 9.2 22.6 30.5 12.1 29.0 37.0
[Experimental Results]: Table 10 indicates that replacing the original failure‑inducing input with the ReduceFix counter‑example improves ChatRepair across every difficulty level. The overall pass@10 rises from 30.5% to 37.0%, an absolute gain of 6.5 percentage points and a relative gain of 21.3%. Looking by difficulty, C-difficulty tasks improve from 41.7% to 45.0% (7.9% relative improvement), D-difficulty tasks from 37.5% to 46.2% (23.2% relative improvement), and the hardest E&F tasks from 10.0% to 16.7% (67.0% relative improvement). Similar lifts appear at pass@5 and pass@1, confirming that the benefit is robust across different sampling $k$. These results demonstrate that reducing the supplied failure-inducing test yields a measurable improvement without requiring any other adjustments to existing APR systems, such as ChatRepair.
[RQ-4] [Findings]: Integrating ReduceFix with ChatRepair increases the overall pass@10 by 21.3% relative to the original configuration. [Insights]: ReduceFix can be adopted as a lightweight add‑on for any other APR system that leverages test cases, providing an immediate boost in repair effectiveness while requiring no changes to the existing APR pipeline.
6. Threats to Validity
Internal validity.
The inherent stochasticity of LLMs may pose threats to internal validity. First, reducer generation is made deterministic by decoding with temperature $=0$; each prompt therefore yields identical code. Second, patch generation retains temperature $=0.8$ so that the model explores diverse fixes. Because developers typically review more than one suggestion, we sample k candidate patches per bug and evaluate pass@1, pass@5, and pass@10. Summarizing results through these statistics converts raw sampling noise into a controlled, reproducible measure, thereby mitigating the threat that randomness alone drives the observed accuracy.
Construct validity.
Improper metrics or biased sampling could distort the study’s reflection of real repair difficulty. First, the compression ratio of ReduceFix often reaches 100%, which hides variation among difficulties. We therefore report both the median and the mean and publish the entire distribution to expose dispersion. Second, pass@k may overstate success when a patch is tuned to a partial test set. Every candidate is thus checked against the full official AtCoder archive, rather than a hand-picked or regenerated subset. Third, dataset bias might arise if only select problems are chosen. We include every AtCoder Beginner Contest problem that satisfies the test size and difficulty level filters within a defined date window, thereby removing any scope for cherry-picking. Together, these steps align the measurements with practical repair goals and mitigate the threats to construct validity identified above.
External validity.
One concern is that ReduceFix might benefit only the pipeline evaluated in this study and fail to transfer to other APR frameworks. To investigate this limitation, we inserted the reducer as a plug‑in into ChatRepair while leaving the rest of its conversation logic unchanged. The modified system achieved a higher pass@10 on the same benchmark, indicating that the method can be integrated into existing APR tools with minimal engineering effort, mitigating the threat of limited applicability.
7. Related Work
7.1. LLM-based Program Repair
Program Repair focuses on automatically or semi-automatically fixing software bugs. It aims to reduce the cost and effort of manual debugging by generating patches that correct faulty behavior in code. LLMs have recently advanced automated program repair, significantly surpassing traditional rule-based or search-based techniques (Yin et al., 2024; Yang et al., 2024; Xia and Zhang, 2022; Yang et al., 2025b, a). ChatRepair is the first work that leverages detailed feedback for each and every patch validated for conversational APR (Xia and Zhang, 2024). Following it, many LLM-based repair pipelines embed the failing test case in the prompt to supply bug context (Yang et al., 2024; Kong et al., 2024; Tang et al., 2024). However, when the failure-inducing test case is long, the LLMs’ context window is exceeded and attention diffuses, producing the “lost-in-the-middle” effect and lowering patch accuracy (Yang et al., 2024; Tian et al., 2023). Reducing the failure-inducing input while preserving the failure is therefore crucial for efficient, focused repair, especially when the goal is to fine-tune compact LLMs on highly informative examples. However, no existing work studies test input reduction in the context of APR pipelines. To our knowledge, our work ReduceFix is the first approach that leverages test input reduction techniques into LLM-based APR.
7.2. Test Input Reduction
Reducing test inputs that trigger bugs is crucial for efficient debugging. Delta debugging is the most popular approach for this purpose in software engineering. Zeller and Hildebrandt initially proposed the first delta debugging algorithm, named ddmin (Zeller, 1999; Zeller and Hildebrandt, 2002). Following ddmin, HDD (Misherghi and Su, 2006) and Perses (Sun et al., 2018) utilize the syntactical structure of the test input to further improve the reduction process. ProbDD (Wang et al., 2021) introduces a probabilistic model to improve ddmin, by estimating the probability of each element being kept in the produced result and prioritizing the reduction of those with high probabilities. GReduce (Ren et al., 2025) assumes that the given input is generated by a test generator and applies ddmin to the generator’s execution trace to reduce the test input. Besides the above delta debugging and its derived algorithms, many domain-specific approaches have been proposed for test input reduction on various domains. For example, CReduce (Regehr et al., 2012) reduces C/C++ programs by iteratively applying pre-defined, well-crafted program transformation rules. ddSMT (Niemetz and Biere, 2013) is proposed for conducting delta debugging on SMT formulas. Binary Reduction (Kalhauge and Palsberg, 2019) is proposed to reduce Java bytecode.
Performing reduction on various types of input formats, such as those in our benchmark tasks, requires manually customizing delta debugging algorithms or designing domain-specific approaches, which is time-consuming and requires domain expertise. Our work addresses this bottleneck by prompting an LLM to automatically generate the input reducer, enabling task-agnostic input minimization that feeds reduced tests directly back to the repair LLMs.
8. Conclusion
Our study addresses two persistent gaps that exist between test-case reduction and LLM-driven automated program repair. First, we demonstrate that LLMs can generate a task-specific reducer from a single prompt, thereby freeing developers from the need for manual customization to accommodate heterogeneous input formats. Second, we place this reduction step within the LLM-based repair loop so that the trimmed failing case provides a precise, high-signal prompt for patch generation. Experiments on the newly built LFTBench benchmark confirm the value of ReduceFix: inputs shrink by up to 100%, and repair success climbs by up to 53.8%. Ablation study shows that the benefit of ReduceFix stems from the combination of brevity and complete failure evidence rather than length trimming alone. An extension experiment further shows that replacing the test input in ChatRepair with the reduced one increases its pass@10 by 21.3%. These improvements demonstrate that integrating input reduction with LLM-based APR materially advances repair effectiveness.
Fully automated reduction unlocks practical applications: programming courses can return minimal counterexamples with fixes to accelerate students’ learning, and continuous-integration pipelines can store smaller regression tests and pinpoint faults for developers more efficiently. Because the reduced input is a plug‑in component, future work can extend ReduceFix into other APR frameworks and even broadly long‑context LLM tasks that profit from concise but information‑rich prompts.